OpenStudy (anonymous):

The equation of a line is y = 2x + 3. What is the equation of the line that is parallel to the first line and passes through (2, –1)?

OpenStudy (anonymous):

the slope is the same: 2 solve: -1 = 2 (2) + b b = -5 equation: y = 2x -5

OpenStudy (anonymous):

2x + y = 6 y = –2x + 3 y = 1/2x + 6 y = –2x – 6

OpenStudy (anonymous):

what are those equations?

OpenStudy (anonymous):

Answers.

OpenStudy (anonymous):

they're incorrect. there can only be one equation

OpenStudy (anonymous):

if you plug (2,-1) into the equations, you be able to see if they are right or not.

OpenStudy (anonymous):

and parallel lines always have the same slope. since the question asks for an equation of a line parallel to the first equation, the slope must be the same which is 2 in this case